a particle moving in a circle with a constant speed covers an angle of 90degrees .the change in its velocity is ??????in terms of v

Grade:11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Aditya

In 90 degree it will cover one fourth of the perimeter

suppose speed is v then time taken t=2∏r/4v=∏r/2v

velocity=displacement/time

displacement=√r2+r2 =r√2

velocity=r√2/t = 2√2 v/∏

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

debdeep bhandari
13 years ago

initial velocity is vi and final velocity is vj,,,,,so the net velocity is vj-vi,,,,whose magnitude is(v2+v2).5